OpenStudy (anonymous):

What is the solution to the equation 14^x + 3 = 21 ?

OpenStudy (anonymous):

so we know 14^x = 18 then use logarithm to find for x so xln14 = ln18 x = ln18/ln14 = ln(18-14) = ln4 if you are in grade 12 or lower ln is like log same function :)